# How Muon’s Spectral Design Benefits Generalization: A Study on Imbalanced Data

### v1

@@ D.7 Effect of Loss Re-weighting
Figure 28: Comparison of GD, Adam, Shampoo and Muon when training a CNN on the Barcoded MNIST dataset from Kunstner et al., (2024), a variant of MNIST with heavy-tailed class imbalance. GD only drives the loss on majority classes towards 0 and makes little progress on the minority classes. In contrast, Adam, Shampoo and Muon drive the loss on both majority and minority classes towards 0.
D.8 MNIST with Heavy-tailed Class Imbalance
@@
Dataset. In this section, we consider the Barcoded MNIST dataset, which is a variant of MNIST with heavy-tailed class imbalance introduced in Kunstner et al., (2024). Barcoded MNIST contains two types of classes: $10$ classes with $5000$ samples each from the original MNIST dataset (majority), and $10\times 2^{10}$ additional classes with $5$ samples each (minority). The images in the minority classes are generated by taking images from the original MNIST dataset, and encoding a $10$-bit barcode into the top left corner of the images, for each of the $10$ original classes.
@@
Model Architecture and Training. Following Kunstner et al., (2024), we train a $2$-layer CNN on this dataset in the full-batch setting. In Fig.˜28, we compare the total train loss as well as train loss on the majority and minority classes (each comprising about $\approx 50\%$ of the total samples) for the four optimizers: GD with momentum, Adam, Shampoo and Muon (we use learning rates $0.005$, $10^{-4}$, $10^{-4}$, and $0.005$, respectively). We use default momentum parameter values for all optimizers. For Shampoo, we use AdamGraftingConfig, and set the stabilization parameter to $10^{-8}$.
@@
Results. Consistent with the results in Kunstner et al., (2024), we observe while that GD only minimizes the loss on the majority classes, and makes negligible progress on the minority classes, Adam minimizes loss on both majority and minority classes. In addition, we find that training with Muon or Shampoo leads to a similar behaviour as Adam: they also minimizes the loss on both majority and minority classes, in contrast to GD.
@@ Appendix E Related Work
Spectrum-aware and Other Adaptive Optimizers. Our work is situated within the growing body of research on spectrum-aware and matrix-level optimizers, which have demonstrated significant empirical success. The development of methods like Shampoo (Gupta et al.,, 2018) and, more recently, Muon (Pethick et al.,, 2025; Jordan et al.,, 2024) has highlighted the potential of optimizers that operate with matrices at the layer level (Carlson et al.,, 2015; Large et al.,, 2024), rather than on vectorized parameters. Recent theoretical inquiries have sought to understand the mechanisms behind their performance. One line of work has focused on characterizing the implicit bias of their canonical form, SpecGD. For instance, Fan et al., (2025); Tsilivis et al., (2024) have shown that SpecGD converges to a max-margin classifier with respect to the spectral norm, though this analysis primarily describes the terminal phase of training. Concurrently, other research has investigated the advantages of adaptive methods like Adam, particularly in settings with heavy-tailed class imbalance (Kunstner et al.,, 2024) or from the perspective of feature learning (Vasudeva et al., 2025b, ). While these works highlight Adam’s advantages on different forms of imbalanced data through various mechanisms, we use this testbed to reveal a fundamentally different principle behind SpecGD’s success: its balanced learning of the data’s principal components, which correspond directly to majority and minority groups.
@@
Further, given that our experiments benchmark spectral methods against strong adaptive baselines, it is worth noting that recent large-scale empirical studies on language model pretraining confirm that well-tuned Adam variants remain highly competitive. Extensive benchmarks have shown that AdamW is a robust and strong baseline (Wen et al.,, 2025; Semenov et al.,, 2025), which is consistent with our experiments in Sec.˜4, where Adam is on par with or sometimes beats Shampoo and Muon. Orvieto and Gower, (2025) showed, via extensive hyperparameter tuning, that $\beta_{1}=\beta_{2}$ is a near-optimal choice for momentum parameters in Adam, contrary to conventional wisdom. They also showed that Signum recovers much of the performance gap between SignGD and Adam, a trend we also observe in our experiments (Fig.˜14). A key factor in its performance relative to SGD is the batch size, as the Adam-SGD gap is most pronounced in large-batch training regimes (Srećković et al.,, 2025; Marek et al.,, 2025). In line with these works, we tune hyperparameters such as the learning rate and momentum parameters for all our experiments (see App.˜D).
@@
Training Dynamics. Our theoretical analysis of training dynamics builds directly upon the framework established by Saxe et al., (2013) and subsequently extended by Gidel et al., (2019). These works provided exact, closed-form solutions for the learning dynamics of deep linear networks trained with standard GD. Their key insight was to analyze the system under the condition that the input and output moment matrices are jointly diagonalizable, which simplifies the complex, non-linear dynamics into a set of decoupled scalar equations. While their analysis revealed how GD prioritizes learning the dominant spectral components of the data, our work adopts this lens and extends it to SpecGD. This extension allows for a direct, analytical comparison between the training trajectories of GD and SpecGD, enabling us to precisely characterize the latter’s more balanced learning rate across all spectral components. Furthermore, our analysis of the bilinear model (Sec.˜3.3) also connects to the literature on the Unconstrained Features Model (UFM), which is often used as a theoretical proxy to study the geometry of feature learning in deep networks, positing a model with enough capacity to freely optimize both the learned representations and the final classification layer (Yang et al.,, 2018; Mixon et al.,, 2020).
@@
Spectral Bias in Deep Learning. More broadly, our work connects to the well-documented phenomenon of spectral bias or simplicity bias in deep learning, where networks trained with (S)GD show a predisposition to learn simpler patterns first. Prior works have demonstrated that networks prioritize learning low-frequency functions before high-frequency ones (Rahaman et al.,, 2019; John Xu et al.,, 2020), and that SGD guides networks to learn functions of progressively increasing complexity (Nakkiran et al.,, 2019; Cao et al.,, 2020). This principle extends even to modern architectures, with recent work showing that Transformers also exhibit a bias towards learning low-sensitivity functions (Bhattamishra et al.,, 2023; Vasudeva et al., 2025a, ). However, this inherent bias towards simplicity can be detrimental, as it can lead models to rely on simple but spurious correlations in the data, which hinders generalization on underrepresented groups (Shah et al.,, 2020; Vasudeva et al., 2025b, ). Our analysis complements this line of work by showing how spectrum-aware optimizers offer a mechanism to somewhat mitigate this bias by promoting more uniform learning across the spectral components of the data.

### v2

@@ D.8 Effect of Loss Re-weighting
D.9 MNIST with Heavy-tailed Class Imbalance
@@
Dataset. In this section, we consider the Barcoded MNIST dataset, which is a variant of MNIST with heavy-tailed class imbalance introduced in kunstner2024heavy. Barcoded MNIST contains two types of classes: $10$ classes with $5000$ samples each from the original MNIST dataset (majority), and $10\times 2^{10}$ additional classes with $5$ samples each (minority). The images in the minority classes are generated by taking images from the original MNIST dataset, and encoding a $10$-bit barcode into the top left corner of the images, for each of the $10$ original classes.
@@
Model Architecture and Training. Following kunstner2024heavy, we train a $2$-layer CNN on this dataset in the full-batch setting. In Fig.˜30, we compare the total train loss as well as train loss on the majority and minority classes (each comprising about $\approx 50\%$ of the total samples) for the four optimizers: GD with momentum, Adam, Shampoo and Muon (we use learning rates $0.005$, $10^{-4}$, $10^{-4}$, and $0.005$, respectively). We use default momentum parameter values for all optimizers. For Shampoo, we use AdamGraftingConfig, and set the stabilization parameter to $10^{-8}$.
@@
Results. Consistent with the results in kunstner2024heavy, we observe while that GD only minimizes the loss on the majority classes, and makes negligible progress on the minority classes, Adam minimizes loss on both majority and minority classes. In addition, we find that training with Muon or Shampoo leads to a similar behaviour as Adam: they also minimizes the loss on both majority and minority classes, in contrast to GD.
D.10 Spectral Analysis for Non-linear Models
In this section, we extend our analysis on the rate of learning of different spectral components beyond linear models. Specifically, we consider two settings with step imbalance ($k/2$ majority and $k/2$ minority classes) trained using cross-entropy (CE) loss: i) one-hidden-layer MLP (with fixed outer layer, as used in Fig.˜1) trained on the data model (DM) with an imbalance ratio of $10$, and ii) a ResNet-18 trained on CIFAR-10 with an imbalance ratio of $20$ (same setting as Sec.˜D.2 for Fig.˜3). In both settings, we track the singular values of the $k\times n$ logit matrix using all $n$ train samples across the $k$ classes as training progresses. Specifically, we quantify the extent of balanced learning by measuring the KL divergence between the uniform probability vector $\tfrac{1}{k}\mathbf{1}$ and the vector of normalized singular values $\tfrac{\operatorname{diag}(\Sigma_{t})}{\operatorname{tr}(\Sigma_{t})}$ over training time, where $\Sigma_{t}$ is the singular value matrix at time $t$.
Figure 31: Comparison of NGD, SignGD, and SpecGD on a one-hidden ReLU MLP on Gaussian mixture data generated using the data model Eq.˜DM. From left to right: train loss, minority- and majority-class test accuracies, and KL divergence between the normalized all-ones vector and the vector with normalized singular values of the logit matrix over time. SpecGD maintains a consistently low KL divergence throughout training. In contrast, NGD and SignGD exhibit an increase in this metric, indicating a period of spectral imbalance.
Gaussian Mixture Setting.
We consider $n=1000,d=200$. Similar to Fig.˜4, we compare optimizers NGD, SignGD and SpecGD in this case, with learning rates $8\times 10^{-2},8\times 10^{-4}$, and $2\times 10^{-2}$, respectively. We use the full batch for training, and use a held out set of $2000$ test samples for reporting test accuracies. The results are reported after averaging over $5$ independent seeds for initialization and data selection. As shown in Fig.˜31, for NGD and SignGD, the KL divergence-based metric initially increases as training progresses, indicating more imbalanced singular values, and becomes smaller in later stages of training, as both majority and minority classes are learned. In contrast, SpecGD maintains a consistently lower value for this metric throughout training, demonstrating that its normalized singular values remain closer to uniform and that spectral components are learned at a more balanced rate. These findings are supported by the corresponding majority- and minority-class accuracies.
Figure 32: Comparison of NMD, Signum, and Muon on imbalanced CIFAR-10 showing the KL divergence between the uniform probability vector and the vector of normalized singular values of the logit matrix. Muon maintains a lower KL divergence throughout training compared to NMD and Signum indicating more balanced learning of the spectral components.
CIFAR-10 with Class Imbalance.
We repeat the experiment described above on the imbalanced CIFAR-10 dataset (same setting as Sec.˜D.2 for Fig.˜3). In Fig.˜32, we observe that Muon similarly demonstrates a more balanced learning of the spectral components. This is supported by the dynamics of majority- and minority-class accuracies (see Fig.˜3).
@@ Appendix E Related Work
Spectrum-aware and Other Adaptive Optimizers. Our work is situated within the growing body of research on spectrum-aware and matrix-level optimizers, which have demonstrated significant empirical success. The development of methods like Shampoo (gupta2018shampoo) and, more recently, Muon (pethick2025trainingdeeplearningmodels; jordan2024muon) has highlighted the potential of optimizers that operate with matrices at the layer level (Carlson2015PreconditionedSD; large2024scalable), rather than on vectorized parameters. Recent theoretical inquiries have sought to understand the mechanisms behind their performance. One line of work has focused on characterizing the implicit bias of their canonical form, SpecGD. For instance, Chen_NSD; tsilivis2024flavors have shown that SpecGD converges to a max-margin classifier with respect to the spectral norm, though this analysis primarily describes the terminal phase of training. Concurrently, other research has investigated the advantages of adaptive methods like Adam, particularly in settings with heavy-tailed class imbalance (kunstner2024heavy) or from the perspective of feature learning (vasudeva2025richsimpleimplicitbias). While these works highlight Adam’s advantages on different forms of imbalanced data through various mechanisms, we use this testbed to reveal a fundamentally different principle behind SpecGD’s success: its balanced learning of the data’s principal components, which correspond directly to majority and minority groups.
@@
Further, given that our experiments benchmark spectral methods against strong adaptive baselines, it is worth noting that recent large-scale empirical studies on language model pretraining confirm that well-tuned Adam variants remain highly competitive. Extensive benchmarks have shown that AdamW is a robust and strong baseline (wen2025fantasticpretrainingoptimizers; semenov2025benchmarkingoptimizerslargelanguage), which is consistent with our experiments in Sec.˜4, where Adam is on par with or sometimes beats Shampoo and Muon. orvieto2025searchadamssecretsauce showed, via extensive hyperparameter tuning, that $\beta_{1}=\beta_{2}$ is a near-optimal choice for momentum parameters in Adam, contrary to conventional wisdom. They also showed that Signum recovers much of the performance gap between SignGD and Adam, a trend we also observe in our experiments (Fig.˜15). A key factor in its performance relative to SGD is the batch size, as the Adam-SGD gap is most pronounced in large-batch training regimes (srećković2025batchsizeproblemrevisiting; marek2025smallbatchsizetraining). In line with these works, we tune hyperparameters such as the learning rate and momentum parameters for all our experiments (see App.˜D).
@@
Training Dynamics. Our theoretical analysis of training dynamics builds directly upon the framework established by saxe2013exact and subsequently extended by gidel2019implicit. These works provided exact, closed-form solutions for the learning dynamics of deep linear networks trained with standard GD. Their key insight was to analyze the system under the condition that the input and output moment matrices are jointly diagonalizable, which simplifies the complex, non-linear dynamics into a set of decoupled scalar equations. While their analysis revealed how GD prioritizes learning the dominant spectral components of the data, our work adopts this lens and extends it to SpecGD. This extension allows for a direct, analytical comparison between the training trajectories of GD and SpecGD, enabling us to precisely characterize the latter’s more balanced learning rate across all spectral components. Furthermore, our analysis of the bilinear model (Sec.˜3.3) also connects to the literature on the Unconstrained Features Model (UFM), which is often used as a theoretical proxy to study the geometry of feature learning in deep networks, positing a model with enough capacity to freely optimize both the learned representations and the final classification layer (yang2018breakingsoftmaxbottleneckhighrank; mixon2020neural).
@@
Spectral Bias in Deep Learning. More broadly, our work connects to the well-documented phenomenon of spectral bias or simplicity bias in deep learning, where networks trained with (S)GD show a predisposition to learn simpler patterns first. Prior works have demonstrated that networks prioritize learning low-frequency functions before high-frequency ones (rahaman2019spectralbiasneuralnetworks; xu2020frequencyprinciple), and that SGD guides networks to learn functions of progressively increasing complexity (nakkiran2019sgdneuralnetworkslearns; cao2020understandingspectralbiasdeep). This principle extends even to modern architectures, with recent work showing that Transformers also exhibit a bias towards learning low-sensitivity functions (bhattamishra2023simplicitybiastransformersability; vasudeva2025transformerslearnlowsensitivity). However, this inherent bias towards simplicity can be detrimental, as it can lead models to rely on simple but spurious correlations in the data, which hinders generalization on underrepresented groups (shah2020pitfalls; vasudeva2025richsimpleimplicitbias). Our analysis complements this line of work by showing how spectrum-aware optimizers offer a mechanism to somewhat mitigate this bias by promoting more uniform learning across the spectral components of the data.